OpenStudy (yueyue):

Given the two equations below, find the average of a and b. 1a+6b=12 6a+1b=16

OpenStudy (yueyue):

The answer must be a whole number.

OpenStudy (nurali):

1a+6b=12---1 6a+1b=16---2 in equation 1 a=12-6b---3 put in equation 2 6(12-6b)+1b=16 72-36b+1b=16 35b=56 b=8/5 put in equation 3 a=12-6(8/5) a=12/5

OpenStudy (yueyue):

Ok thank you so much!

OpenStudy (nurali):

My Pleasure.